Constants Part A Three point charges are located on the positivex axis of a coordinate system. Charge q1 1.0 nC is 2.0 cm from the origin, charge q2 -2.0 nC is 4.0 cm from the origin and charge q3 4.5 nC located at the origin. What is the net force (a)magnitude and (b) direction) on charge q 1.0 nC exerted by the other two charges? Express your answer using two significant figures You may want to review (Pages 690 - 695) For related problemsolving tips and strategies, you may want to view a Video Tutor Solution of Vector addition of electric forces on a line Submit Previous Answers Request Answer
Answer #1

q1 = 1.0nC located at x = 2.0 cm
q2 = -2.0nC located at x = 4.0 cm
q3 = 4.5nC located at x = 0 cm

(a) Net Force on 1.0 nC charge

From Coulomb's law, net force on 1.0 nC charge is written as:

Fnet = 1/4πϵo * q1*q3/(2*10^-2)2 - 1/4πϵo * q1*q2/(2*10^-2)^2

Fnet = 1/4πϵo * 1*10^-9 *(4.5*10^-9+2*10^-9)/4*10^-4

Fnet =9*10^9 * (1*10^-9*6.5*10^-9/4*10^-9)

= 14.625N

(b) Direction of force on 1.0 nC Charge

In the calculation of force in part(a), the value of force is positive,
which indicates that the direction of force is towards +x